OpenStudy (anonymous):

The value of a car x years from now is given by v(x) = 2000(0.75)x. What is the annual compound interest rate of depreciation? A. 15% B. 25% C. 75% D. 105%

OpenStudy (anonymous):

@ninjawolfjj

OpenStudy (anonymous):

I got B for my answer

OpenStudy (anonymous):

Yeah its B, because 2000 * 0.75 = 1500 and 2000 - 1500 = 500 500+500+500+500 = 2000

OpenStudy (anonymous):

And that makes 500 25%

OpenStudy (anonymous):

GTG bye

OpenStudy (anonymous):

aww man i had a few more questions, k bye:)